OpenStudy (anonymous):

13/36=0.361111111 how would i reverse a decimal back to a fraction

OpenStudy (amoodarya):

a=0.36111111111111... 100a =36.11111111111.... 1000a =361.1111111111.... subtract 1000a-100a=361.1111111...-36.111111...=325 so a=325/900 then simplify that

OpenStudy (anonymous):

why is it over 900 though i dont understand

OpenStudy (anonymous):

because im trying to get 0.36111111111... back to 13/36

OpenStudy (unklerhaukus):

\[x=0.361111111\] \[9x=0.361111111\times 9\]\[=0.3\times9+0.06\times9+0.00\dot1\]\[=2.7+0.54+0.00\dot9\]\[=2.7+0.54+0.01\] \[9x=2.7+0.55\]\[\quad=3.25\]\[x=3.25/9\]\[x=\frac{3+1/4}9\]

OpenStudy (unklerhaukus):

\[x=\frac{13/4}9\]

OpenStudy (amoodarya):

if you simplify the 325/900=13/36

OpenStudy (anonymous):

where did u get the 900 from is what im stuck at

OpenStudy (unklerhaukus):

1000a-100a=900a

OpenStudy (anonymous):

what is the fastest way to simplyfly 325/900

OpenStudy (unklerhaukus):

cancel the common factor of 25